# Evolutionary Optimization of Physics-Informed Neural Networks: Evo-PINN Frontiers and Opportunities

### v3

@@ Abstract
Deep learning models trained on finite data lack a complete understanding of the physical world. On the other hand, physics-informed neural networks (PINNs) are infused with such knowledge through the incorporation of mathematically expressible laws of nature into their training loss function. By complying with physical laws, PINNs provide advantages over purely data-driven models in limited-data regimes and present as a promising route towards Physical AI. This feature has propelled them to the forefront of scientific machine learning, a domain characterized by scarce and costly data. However, the vision of accurate physics-informed learning comes with significant challenges. This work examines PINNs for the first time in terms of model optimization and generalization, shedding light on the need for new algorithmic advances to overcome issues pertaining to the training speed, precision, and generalizability of today’s PINN models. Of particular interest are gradient-free evolutionary algorithms (EAs) for optimizing the uniquely complex loss landscapes arising in PINN training. Methods synergizing gradient descent and EAs for discovering bespoke neural architectures and balancing multiple terms in physics-informed learning objectives are positioned as important avenues for future research. Another exciting track is to cast evolutionary as a meta-learner of generalizable PINN models. To substantiate these proposed avenues, we further highlight results from recent literature to showcase the early success of such approaches in addressing the aforementioned challenges in PINN optimization and generalization.
@@
Physics-informed neural networks, evolutionary optimization, neuroevolution, generalizability, survey.
@@ I Introduction
Figure 1: Google Scholar survey of the literature shows the popularity of PINN over the years between 2019 and 2024, and PINN studies in diverse scientific areas. Data collected in November 2024.
@@
Branch Applications Reference
Fluid Mechanics Turbomachinery Design [7]
Airfoil Design [8]
Solid Mechanics Topology Optimization [9, 10]
Vehicle Crashworthiness Design [11]
Additive Manufacturing [12, 13]
Digital Twin [14, 15]
@@
Quantum & Particle Physics Diffuse Optical Tomography [17]
Nuclear Fusion Physics Analysis [18, 19]
@@
The shortcomings of standard data-driven models have seeded the field of scientific machine learning where minimizing deviations from the known laws of nature serves to boost model precision and data efficiency [21, 22]. Associated methods offer unique pathways to achieving rationalizable AI, defined in [23], that is cognizant of fundamental physics principles. A myriad of approaches to instil physics consistency into learned models can be crafted, with the most suitable methods being naturally dependent on the task at hand and the type of knowledge available [24, 25, 26, 27]. Some instantiations include the design of model architectures that preserve invariances or symmetries representative of the underlying system [28], data augmentation with theory-based synthetic data (e.g., first principles numerical simulations) [29], or introduction of learning biases into the training process [30].
@@
This effort focuses on the growing body of literature on physics-informed neural networks (PINNs), popularized in recent years by the work of Raissi et al. [31]. The essential idea of a PINN—where mathematically-defined learning biases are directly incorporated into a neural network’s loss function—is simultaneously decades old, with early work in the 1990s [32, 33], and also a breath of fresh air for AI. This physics-informed loss is amenable to various forms of scientific knowledge and theories, including conservation laws and/or governing equations in the form of ordinary and partial differential equations (ODEs and PDEs), yet retains immense flexibility in the choice of network architectures (e.g., multilayer perceptron, convolutional or graph neural networks) and learning algorithms (e.g., gradient-based or gradient-free). This makes it possible for scientific knowledge uncovered over centuries, from the micro-scale evolution of quantum-mechanical systems to the macro-scale behavior of solids, fluids and propagation of electromagnetic fields, to be incorporated into the learned models of today.
@@
In theory, PINNs compensate for a lack of labeled data by augmenting deep learning with physics principles. The potential of such physics-informed models has spawned significant interest across various scientific domains. Trends from a Google Scholar survey of the literature over 6 recent years (2019-2024) are depicted in Fig. 1. The figure reveals rapid growth in the popularity of PINNs in different branches of science, with many demonstrations being developed using open software libraries such as JAX [34], NVIDIA Modulus [35] (previously referred to as SimNet [36]), SciANN [37], and DeepXDE [38].
@@
Despite the broad applicability of PINNs (ref. to Table I for a summary of potential applications of PINNs), the switch from standard data-driven loss functions to physics-informed learning objectives has brought unforeseen difficulties in optimizing the latter’s uniquely complex landscape [39, 40, 41, 42, 43]. These training difficulties undermine the theoretical advantages of PINNs, and are currently a barrier to their adoption in place of established numerical methods in science and engineering. Existing surveys [27, 44, 45, 46, 47, 48] review PINN methods, applications, benchmark data, limitations, trends and prospects from interdisciplinary viewpoints, but a common point of emphasis is that efficient optimization of PINNs remain an open question direction [48].
@@
Separately, despite the physics-informed loss, the physical validity of a PINN is limited to scenarios encountered during training, even after a complicated training process. They do not inherently generalize the learned physics to new scenarios beyond their training regime—e.g., changing PDE parameters, initial conditions or boundary conditions—thereby continuing to suffer from generalizability issues similar to standard data-driven learning algorithms.
@@
In light of the above, this paper examines PINNs for the first time through the lens of their optimization and generalization challenges. Notably, most PINN implementations today make use of gradient-based optimization algorithms. The ubiquity of the method may be attributed to inertia from the primacy of stochastic gradient descent (SGD) in conventional data-rich deep learning tasks. However, SGD is no silver bullet. Relative to standard deep learning, physics-informed loss functions are not only highly non-convex but also exhibit disparities in the magnitudes of back-propagated gradient signals between the data and physics loss terms (see II-A). These issues lead to pathologies that hinder effective training [49, 50]. The resulting optimization trajectory is littered with spurious minima, making it extremely challenging to find globally optimum solutions that satisfy physics. There is a high chance for gradient descent to get trapped in a local minimum that fails to satisfy the requisite physics, a possibility discussed in greater detail in Section III.
@@
In order to fulfill the vision of data-efficient scientific machine learning, this paper asserts the need for algorithmic advances beyond gradient descent to move the needle for PINNs. Fig. 2 provides a schematic summary of some of the main computational challenges in this regard, and some corresponding proposed solutions. In Section IV, arguments are presented in support of our main claim: that algorithmic ideas developed within the computational intelligence community may be uniquely well-suited for physics-informed learning. Of particular interest are the gradient-free evolutionary algorithms (EAs) that suffer less from spurious local minima due to their exploration-oriented global search capacity [51, 52]. In addition to optimizing network parameters, EAs enable the discovery of bespoke architectures for physics-informed learning [53]. Evolutionary mechanisms could also prove useful for guiding training data generation in areas with higher errors on governing physics equations, having shown evidence of improved convergence to correct solutions [54].
An associated research direction is the synergy of EAs with gradient-based training (in the spirit of memetic algorithms), thereby marrying the best of both worlds. While population-based EAs may be advantageous in navigating highly non-convex loss landscapes and balancing trade-offs between conflicting terms in physics-informed learning objectives, local search via gradient descent can facilitate rapid convergence to precise solutions. Yet another exciting track is to cast evolutionary as a meta-learner of generalizable PINN models [55]. This may be achieved through a bilevel formulation where an upper-level evolutionary searches for PINN configurations that can be quickly fine-tuned for various downstream physics tasks at the lower-level.
@@
The advancement of Evo-PINNs is expected to have far-reaching impacts on establishing PINN methodologies at the forefront of scientific machine learning. Yet, surprisingly, there are only a handful of related works in the literature that explore such ideas, implying huge scope for future research and development.
@@ II Overview of PINN Methodologies
PINNs are generally designed to learn a mapping between the input and output variables while satisfying specified mathematical constraints that represent the physical phenomenon or dynamical process of interest [31]. Given their ubiquity in modern science and engineering, PDEs are one of the most studied physics priors integrated in PINNs. Without loss of generality, we consider a physics-informed learning problem with input variables $x$ (spatial dimension) and $t$ (time dimension), and an output variable of interest $u$ satisfying PDEs of the general form:
@@
where the general differential operator $\mathcal{N}_{\vartheta}[u(x,t)]$ contains PDE parameters $\vartheta$ and can include linear and/or nonlinear combinations of temporal and spatial derivatives of $u$, and $h(x,t)$ is an arbitrary source term in the domain $x\in\Omega,t\in(0,T]$. Equation 1b specifies the initial condition (IC), $u_{0}(x)$, at time $t$=00, and Equation 1c specifies the boundary condition (BC) such that $\mathcal{B}[u(x,t)]$ equates to $g(x,t)$ at the domain boundary $\partial\Omega$, where $\mathcal{B}[\cdot]$ can either be an identity operator (Dirichlet BC), a differential operator (Neumann BC), or a mixed identity-differential operator (Robin BC).
Figure 3: PINN handles both scientific simulation and scientific discovery. Example applications are: 1) accurate simulation of transient vortex shedding behind a cylinder obtained by learning a PINN for a single problem solely using the governing physics (2D transient N-S equations); and 2) inference of the wave propagation speed, $c$, given sparse observations (locations demarcated by $\boldsymbol{\times}$), by learning an inference PINN model with a partially specified 1D transient wave equation.
@@ II-A Physics-Informed Loss
where the relative weights $\lambda_{physics}$ and $\lambda_{data}$ control the trade-off between the two components. The physics-informed component can consist of multiple PDE, IC, and BC loss terms:
@@
where $f(\cdot;\boldsymbol{w})$ is the PINN output function with network parameters $\boldsymbol{w}$, evaluated on a set of label-free training points (collocation points) sampled from the respective spatio-temporal domain. An illustrative schematic of a PINN model is presented in Fig. 3. Evaluation of differential operators, e.g. $\frac{d\ f(\cdot;\boldsymbol{w})}{dt}$, $\frac{d\ f(\cdot;\boldsymbol{w})}{dx}$ and $\frac{d^{2}f(\cdot;\boldsymbol{w})}{dx^{2}}$, are generally required for computation of the physics-informed loss on these training samples. They can be conveniently computed by automatic differentiation [31] and/or numerical differentiation [56, 57] methods at any sample location.
Fundamentally, PINNs arrive at an accurate and physics-compliant prediction for a desired physics scenario by forcing its outputs $f(\cdot;\boldsymbol{w})$ to satisfy Equation 1 through minimizing the physics-informed loss (Equation 3). The effectiveness of the optimization of network parameters $\boldsymbol{w}$ through the physics-informed loss is a primary determinant of the accuracy of the PINN model. In contrast, the typical data-driven loss (as common in a DNN) computes the MSE between the model prediction $f(x_{i},t_{i};\boldsymbol{w})$ and available target $u_{i}^{label}$ over $i$=$1,...,n$ labeled samples:
@@ II-B Physics-Informed Learning Advantages
The advantages of physics-informed learning are demonstrated in Fig. 4. Given only a few labeled data ($n$=$5$), a DNN model (with sufficient representation capacity) remains unable to correctly reconstruct the entire system, although it fits very well on all training samples. On the other hand, a PINN model can accurately learn the solution from the same set of labeled data. Given that the physics constraints are enforced on a set of label-free training samples ($n_{pde}$=$1000$) uniformly drawn from the computation domain utilizing the aforementioned PINN loss, a PINN can learn the solution even without any labeled data. The governing physics in this example is the convection-diffusion equation $\alpha\frac{du}{dx}-\frac{d^{2}u}{dx^{2}}=0$ with PDE parameter $\alpha$, a ubiquitous physics model in science and engineering that describes the distribution of a scalar quantity (e.g., mass, energy, temperature) in the presence of convective transport and diffusion [58].
@@ II-C PINN for Scientific Simulation
Compared to classical numerical solvers employing the finite difference, finite element, or finite volume method [60], the PINN approach has the significant advantage of being mesh-free. Meshing, which is the process of domain discretization, is an essential, yet non-trivial task. Classical numerical methods can fail to find the correct solution if the meshing is not appropriately done. On the other hand, PINNs can provide a solution without similar mesh requirements. Additionally, the solution obtained through PINN is differentiable, and its compact representation requires lower memory demand for storage. This makes PINNs a promising alternative to traditional numerical methods for simulating physics phenomena; an extensive review is presented in [44].
While most PINN studies address single physics problem, a PINN model can also incorporate additional input parameters, such as geometry, Reynolds number, or material properties [61, 62, 63, 64]. These models, also referred to as parameterized PINN, can then provide predictions across a set of physics scenarios, rather than just one.
@@ II-D PINN for Scientific Discovery
In addition to directly simulating physical behavior, PINNs can be seamlessly extended to various kinds of hypothesis inference problems useful for scientific discovery. These problems typically involve an incomplete and potentially noisy set of observed data points, and the goal is to infer information about the system through optimizing the hypothesis space. The loss function used to train PINN models for inference consists of both physics-informed (to capture relevant physics) and data (observations) components (Equation 2). The hypothesis spaces of the inference problem are included as optimization variables along with the network parameters during training; see example in Fig. 3.
The PINN models are uniquely suited for hypothesis inference when compared to numerical solvers and can effectively synergize theory and observation data from various sources such as sensors and imaging, even for ill-posed PDE problems. It can be combined with probabilistic inference methods (e.g. Bayesian-PINN [65]), with the probabilistic framework offering a mathematical tool to handle uncertainties arising from both noisy data and physics assumptions [65, 66, 67].
In the literature, PINN models have been applied to: (1) Inference of system-level parameters such as the underlying governing equations or constitutive relationships or system parameters (e.g., Reynolds number, wave speed) based on observations [68, 69, 70, 71, 72, 73]. (2) Inference of unobserved field variables across the entire domain based on limited observations, e.g., predicting pressure in a particle imaging velocimetry experiment with only velocity observations [74]. Many have employed PINNs towards the inference of domain field variables in various problems spanning fluid dynamics [75], heat transfer [76] and biomedical sciences [77, 78]. (3) PDE-constrained design optimization towards a set of desired (observation) outcomes. The PINN is used as an adjoint optimizer, enabling the design parameters to be optimized towards the desired outcome while fulfilling all specified relevant physics during training. This has been successfully demonstrated on topology optimization [79], metamaterial [80] and biomimetics [81] design problems.

### v5

@@ Abstract
Deep learning models trained on finite data lack a complete understanding of the physical world. On the other hand, physics-informed neural networks (PINNs) are infused with such knowledge through the incorporation of mathematically expressible laws of nature into their training loss function. By complying with physical laws, PINNs provide advantages over purely data-driven models in limited-data regimes and present as a promising route towards Physical AI. This feature has propelled them to the forefront of scientific machine learning, a domain characterized by scarce and costly data. However, the vision of accurate physics-informed learning comes with significant challenges. This work examines PINNs in terms of model optimization and generalization, shedding light on the need for new algorithmic advances to overcome issues pertaining to the training speed, precision, and generalizability of today’s PINN models. Of particular interest are gradient-free evolutionary algorithms (EAs) for optimizing the uniquely complex loss landscapes arising in PINN training. Methods synergizing gradient descent and EAs for discovering bespoke neural architectures and balancing multiple terms in physics-informed learning objectives are positioned as important avenues for future research. Another exciting track is to cast EAs as a meta-learner of generalizable PINN models. To substantiate these proposed avenues, we further highlight results from recent literature to showcase the early success of such approaches in addressing the aforementioned challenges in PINN optimization and generalization.
@@
Physics-informed neural networks, scientific machine learning, evolutionary optimization, neuroevolution, generalizability.
@@ I Introduction
Figure 1: Google Scholar search of the literature shows the popularity of PINN over the years between 2019 and 2024, and PINN studies in diverse scientific areas. Data collected in November 2024.
@@
Branch Applications
Fluid Mechanics Turbomachinery design [7], airfoil design [8]
Solid Mechanics Topology optimization [9, 10], vehicle crashworthiness design [11], additive manufacturing [12, 13], digital twin [14, 15]
@@
Quantum & Particle Physics Diffuse optical tomography [17], nuclear fusion physics analysis [18, 19]
@@
The shortcomings of standard data-driven models have seeded the field of scientific machine learning where minimizing deviations from the known laws of nature serves to boost model precision and data efficiency [21, 22]. Associated methods offer unique pathways to rationalizable AI, defined in [23], that is cognizant of fundamental physics principles. A myriad of approaches to instil physics consistency into learned models can be crafted, with the most suitable methods being naturally dependent on the task at hand and the type of knowledge available [24, 25, 26, 27]. Some instantiations include the design of model architectures that preserve invariances or symmetries representative of the underlying system [28], data augmentation with theory-based synthetic data (e.g., first principles numerical simulations) [29], or introduction of learning biases into the training process [30].
@@
This effort focuses on the growing body of literature on physics-informed neural networks (PINNs), popularized in recent years by the work of Raissi et al. [31]. The essential idea of a PINN—where mathematically-defined learning biases are directly incorporated into a neural network’s loss function—is simultaneously decades old, with its earlier work in the 1990s [32, 33], and also a breath of fresh air for AI. This physics-informed loss is amenable to various forms of scientific knowledge and theories, including conservation laws and/or governing equations in the form of ordinary and partial differential equations (ODEs and PDEs), yet retains immense flexibility in the choice of network architectures (e.g., multilayer perceptron, convolutional or graph neural networks) and learning algorithms (e.g., gradient-based or gradient-free). This makes it possible for scientific knowledge uncovered over centuries, from the micro-scale evolution of quantum-mechanical systems to the macro-scale behavior of solids, fluids, and propagation of electromagnetic fields, to be incorporated into the learned models of today, thereby compensating for a lack of labeled data.
@@
The potential of such physics-informed models has spawned significant interest across various scientific domains. Trends from a Google Scholar search of the literature in the last six years (2019-2024) are depicted in Fig. 1. The figure reveals rapid growth in the popularity of PINNs in different branches of science, with many demonstrations being developed using open software libraries such as JAX [34], NVIDIA PhysicsNeMo  [35], SciANN [36], and DeepXDE [37].
@@
Despite the broad applicability of PINNs (ref. to Table I for some potential applications), the switch from standard data-driven loss functions to physics-informed learning objectives has brought unforeseen difficulties in optimizing the latter’s uniquely complex landscape [38, 39, 40]. These training difficulties undermine the theoretical advantages of PINNs, and are currently a barrier to their adoption in place of established numerical methods in science and engineering. Prior surveys [27, 41, 42, 43, 44, 45] review PINN methods, applications, limitations, trends, and prospects from interdisciplinary viewpoints, but a common point of emphasis is that efficient optimization of PINNs remains an open problem [44].
@@
Separately, despite the physics-informed loss, the physical validity of a PINN is limited to scenarios encountered during training, even after a complicated training process. Most PINNs are designed to model a specific PDE instance, and model training can be highly sensitive to choice of architecture and optimization hyperparameters. Consequently, reusing identical hyperparameter configurations to train new PINN models within the same PDE family often fails to produce satisfactory results. They do not inherently generalize learned physics to new scenarios beyond their training regime—e.g., changing PDE parameters, initial conditions or boundary conditions—thereby continuing to suffer from generalizability issues similar to standard data-driven learning algorithms.
@@
In light of the above, this paper examines PINNs through the lens of their optimization and generalization challenges. As evidenced by Fig. 1, there is a vast amount of innovative PINN literature, however, this paper specifically focuses on a curated subset that best illustrates both these optimization and generalization challenges, and the exciting opportunities for novel, targeted algorithms and methods. Notably, most PINN implementations today make use of gradient-based optimization algorithms. The ubiquity of the method may be attributed to inertia from the primacy of stochastic gradient descent (SGD) in conventional data-rich deep learning tasks. However, SGD is no silver bullet. Relative to standard deep learning, physics-informed loss functions are not only highly non-convex but also exhibit disparities in the magnitudes of back-propagated gradient signals between the data and physics loss terms (see III-A1). These issues lead to pathologies that hinder effective training [46, 47]. The resulting optimization trajectory is littered with spurious minima, making it extremely challenging to find globally optimum solutions that satisfy physics. There is a high chance for gradient descent to get trapped in a local minimum that fails to satisfy the requisite physics, a possibility discussed in greater detail in Section III.
@@
Fig. 2 provides a schematic summary of some of the main computational challenges inhibiting the vision of physics-informed machine learning towards AI for Science, and some corresponding proposed solutions. In Section IV, arguments are presented in support of our main claim: that algorithmic ideas developed over the years within the computational intelligence community may be uniquely well-suited for physics-informed learning. Of particular interest are the gradient-free evolutionary algorithms (EAs) that suffer less from spurious local minima due to their exploration-oriented global search capacity [48, 49]. An associated research direction is the synergy of EAs with gradient descent (in the spirit of memetic algorithms), thereby marrying the best of both worlds. While population-based EAs may be advantageous in navigating highly non-convex loss landscapes and balancing trade-offs between conflicting objectives, local search via gradient descent can facilitate rapid convergence to precise solutions.
In addition to optimizing network parameters, EAs enable the discovery of bespoke architectures for physics-informed learning [50]. Evolutionary mechanisms could also prove useful for guiding training data generation in areas with higher errors on governing physics equations, having shown evidence of improved convergence to correct solutions [51]. Accurate solutions to novel PDEs can be rapidly obtained through evolutionary transfer optimization, leveraging multiple knowledge sources during optimization for specific target tasks [48]. Another exciting track is to cast EAs as a meta-learner of generalizable PINN models [52]. This may be achieved through a bilevel formulation where an upper-level EA searches for PINN configurations that can be quickly fine-tuned for various downstream physics tasks at the lower-level.
@@
The advancement of Evo-PINNs is expected to have far-reaching impact on establishing PINN methodologies at the forefront of scientific machine learning. However, there are only a handful of related works in the literature, implying huge scope for future research and development.
@@ II Overview of PINN Methodologies
PINNs are designed to learn a mapping between the input and output variables while satisfying specified mathematical constraints that represent the physical phenomenon or dynamical process of interest [31]. Given their ubiquity in modern science and engineering, PDEs are one of the most studied physics priors integrated in PINNs. Without loss of generality, we consider a physics-informed learning problem with input variables $x$ (spatial dimension) and $t$ (time dimension), and an output variable $u$ satisfying PDEs of the general form:
@@
where the general differential operator $\mathcal{N}_{\vartheta}[u(x,t)]$ contains PDE parameters $\vartheta$ and can include linear and/or nonlinear combinations of temporal and spatial derivatives of $u$, and $h(x,t)$ is an arbitrary source term in the domain $x\in\Omega,t\in(0,T]$. Equation 1b specifies the initial condition (IC), $u_{0}(x)$, at time tt=0, and Equation 1c specifies the boundary condition (BC) such that $\mathcal{B}[u(x,t)]$ equates to $g(x,t)$ at the domain boundary $\partial\Omega$, where $\mathcal{B}[\cdot]$ can either be an identity operator (Dirichlet BC), a differential operator (Neumann BC), or a mixed identity-differential operator (Robin BC).
[…]
@@ II-A Physics-Informed Loss
where the relative weights $\lambda_{physics}$ and $\lambda_{data}$ control the component-wise trade-off. The physics-informed component can consist of multiple PDE, IC, and BC loss terms:
@@
where $f(\cdot;\boldsymbol{w})$ is the PINN output function with network parameters $\boldsymbol{w}$, evaluated on a set of label-free training points (collocation points) sampled from the respective spatio-temporal domain. An illustrative schematic of a PINN model is presented in Fig. 3. Evaluation of differential operators, e.g., $\frac{d\ f(\cdot;\boldsymbol{w})}{dt}$, $\frac{d\ f(\cdot;\boldsymbol{w})}{dx}$, and $\frac{d^{2}f(\cdot;\boldsymbol{w})}{dx^{2}}$, are generally required for computation of the physics-informed loss on these training samples. They can be conveniently computed by automatic differentiation [31] and/or numerical differentiation [53, 54] methods at any sample location.
Fundamentally, PINNs arrive at an accurate and physics-compliant prediction for a desired physics scenario by forcing its outputs $f(\cdot;\boldsymbol{w})$ to satisfy Equation 1 through minimizing the physics-informed loss (Equation 3). The effectiveness of the optimization of network parameters $\boldsymbol{w}$ through the physics-informed loss is a primary determinant of the accuracy of the PINN model. In contrast, the typical data-driven loss (as common in a DNN) computes the MSE between the model prediction $f(x_{i},t_{i};\boldsymbol{w})$ and available target $u_{i}^{label}$ over ii=1,,n1,...,n labeled samples:
@@ II-B Physics-Informed Learning Advantages
The advantages of physics-informed learning are demonstrated in Fig. 4. Given only a few labeled data (nn=55), a DNN model (with sufficient representation capacity) remains unable to correctly reconstruct the entire system, although it fits all training samples well. In contrast, a PINN model can accurately learn the solution from the same set of labeled data. By enforcing physics constraints on a set of label-free training samples ($n_{pde}$=10001000) uniformly drawn from the computation domain utilizing the aforementioned PINN loss, a PINN can learn the solution even without any labeled data. The governing physics in this example is the convection-diffusion equation $\alpha\frac{du}{dx}-\frac{d^{2}u}{dx^{2}}=0$ with PDE parameter $\alpha$, a ubiquitous physics model that describes the distribution of a scalar quantity (e.g., mass, energy, temperature) in the presence of convective transport and diffusion [55].
@@ II-C PINN for Scientific Simulation
Compared to classical numerical solvers employing the finite difference, finite element, or finite volume method [57], PINNs have the significant advantage of being mesh-free. Meshing, which is the process of domain discretization, is an essential, yet non-trivial task in classical numerical methods. Numerical methods can fail to find the correct solution if the meshing is not appropriately done. On the other hand, PINNs can provide a solution without similar mesh requirements. Additionally, the solution obtained through PINN is differentiable, and its compact representation requires lower memory demand for storage. This makes PINNs a promising alternative to traditional numerical methods for simulating physics phenomena; an extensive review is presented in [41].
While most PINN studies address single physics problem, a PINN model can also incorporate additional input parameters, such as geometry, Reynolds number, or material properties [58, 59, 60]. These models, also referred to as parameterized PINN, can then provide predictions across a set of physics scenarios, rather than just one.
@@ II-D PINN for Scientific Discovery
In addition to directly simulating physical behavior, PINNs can be seamlessly extended to various hypothesis inference problems common in scientific discovery. These problems typically involve an incomplete and potentially noisy set of observed data points, and the goal is to infer information about the system through optimizing the hypothesis space. The loss function used consists of both physics-informed (to capture relevant physics) and data (observations) components (Equation 2). The hypothesis spaces of the inference problem are included as optimization variables along with the network parameters during training; see example in Fig. 3.
The PINN models are uniquely suited for hypothesis inference when compared to numerical solvers and can effectively synergize theory and observation data from various sources such as sensors and imaging, even for ill-posed PDE problems. It can be combined with probabilistic inference methods (e.g., Bayesian-PINN [61]), with the probabilistic framework offering a mathematical tool to handle uncertainties arising from both noisy data and physics assumptions [61, 62].
In current literature, PINN models have been applied to: 1.) Inference of system-level parameters such as the underlying governing equations, constitutive relationships or system parameters (e.g., Reynolds number, wave speed) based on observations [63, 64, 65, 66, 67]. 2.) Inference of unobserved field variables across the entire domain based on limited observations, e.g., predicting pressure from velocity in particle imaging velocimetry [68]. Many have employed PINNs towards the inference of domain field variables in various problems spanning fluid dynamics [69], heat transfer [70], and biomedical sciences [71, 72]. 3.) PDE-constrained design optimization towards a set of desired (observation) outcomes. The PINN is used as an adjoint optimizer, enabling the design parameters to be optimized towards the desired outcome while fulfilling all specified relevant physics during training. This has been successfully demonstrated on topology optimization [73] and metamaterial [74] design problems.